Use the following exchange quotes to help answer problems 18-19 NYSE NASDAG Bid Bid Ask 150.13 Company Y 18. Currently you have $120.00 in cash and 1000 shares of Stock Y in your brokerage account. You plan to sell your 1000 shares of Stock Y and use the proceeds to purchase as many shares of Stock X as possible. Assume you cannot borrow or buy partial shares of stock; you will also buy or sell at the best price offered. Based on the quotes provided, the cash balance in your brokerage account after you purchase the shares of Stock X is closest to: A. $160.89) B. $59.11 C $87.95 D. $89.24 E. $106.10 19. Based on the bid-ask spreads for Stock X & Y, theNASDAQǐsmoreligvidthantheNYSE whileStockY1s more liquid than Stock X. Evaluate the underlined words in italies. True or False? Hint: Calculate spread as a percentage of the midpoint of the bid-ask quotes. A. True B. False 20. During lecture we compared Warren Buffets Berkshire Hathaway (BRK.A), which has a market cap of around $480B, to Amazon (AMZN), which has a market cap of around $790B. larger the market cap of a stock, the greater the stocks liquidity. Evaluate the underlined words in italics. True or False? A. True B. False 21. During lecture we compared the stock market to the bond market. Which of the following statements (1- IlI) is (are) most likely FALSE I Stocks are generally more liquid than bonds Il. There is more money invested in the US stock market than the US bond market Il The fact that bonds are priced relatively infrequently made things worse during the housing crisis С. ії, 111 E. None of the statements
Answer #1

18)

A B C D E F G H I J
2
3 NYSE NASDAQ
4 Bid Ask Bid Ask
5 Company X $149.55 $150.45 $149.87 $150.13
6 Company Y $127.88 $128.12 $127.70 $128.30
7
8
9 Cash Balance $120
10 Number of share of Stock Y 1000
11
12 Bid price is the price at which stock can be sold and ask price is the price at which stock can be bought.
13
14 Thus stock Y will be sold at the highest Bid price.
15 Selling Price of Stock Y $127.88 <-- =MAX(D6,F6)
16
17 Proceed from Sale of stock Y $127,880.00 <-- =D10*D15
18
19 Total Amount available to buy stock X $128,000.00 <-- =D9+D17
20
21 Stock X will be bought at the lowest ask price.
22 Buying Price of stock $150.13 <-- =MIN(E5,G5)
23
24 Total number of shares to be bought 852.5944182 <-- =D19/D22
25 Number of shares X to be bought 852
26
27 Amount Required to buy the shares   $127,910.76 <-- =D25*D22
28
29 Balance in the brokerage account $89.24 <-- =D19-D27
30
31 Hence amount available in brokerage account $89.24
32 Thus the option (D) is correct.
33

F